OpenStudy (kc_kennylau):

Find the derivative of \(2x+3y^2+3x^2y=4x^2\)

OpenStudy (tkhunny):

\(2 + 6yy' + \)... Two more terms to go!

OpenStudy (shamil98):

implicit differentiation, fun.

OpenStudy (shamil98):

\[\huge 2x + 3y^2 + 3x^2y = 4x^2\] \[\huge 2 + 6yy' + 6xy + 3x^2y' = 8x\] \[\huge y'(6y+3x^2) = 8x - 6xy - 2\] \[\huge y' = \frac{ 2(4x-3xy-1)}{ 3(x^2 + 2y) }\]

OpenStudy (anonymous):

good job shamil

OpenStudy (kc_kennylau):

Wow thank you @shamil98 :D

OpenStudy (shamil98):

mhm

OpenStudy (tkhunny):

No good @shamil98 Please pay better attention to the Code of Conduct.

OpenStudy (kc_kennylau):

@tkhunny A Cantonese saying: "Rules are dead. Humans are alive." (I translated it so don't google it) @shamil98 knows that I know this basic knowledge and he knows that when he just solves the answer with steps, I'll be able to understand. The rule's purpose is just to make people understand the whole concept instead of just the question, but he knows that this is the best way for me to understand the whole concept. Different people have different kinds of learning method, and I know that guiding will not be the best method for me, while examples is the best method for me. Therefore, I don't mind it.

OpenStudy (tkhunny):

@kc_kennylau In this sort of forum, it is impossible to tell the difference between a useful example and simply doing your homework for you. Thus, to avoid the error of doing someone else's homework, just don't do it. Thus the Code of Conduct. No one learns mathematics better by following someone else's work compared to how much one will learn by thinking it through an doing the work one's self. I do not believe it is acceptable simply to discard rules because of your belief system or because of a saying from a background that isn't even yours. There may be other ways to get along, but generally, I would suggest that if your beliefs are not consistent with the established rules, feel free to take your beliefs elsewhere. Note: I represent only myself and absolutely not OpenStudy or its moderators, owners, or sponsors. My opinions are my own - sometimes they are welcome.

OpenStudy (kc_kennylau):

"a saying from a background that isn't even yours" What do you mean? I was born in HK, raised up in HK and am living in HK. What do you mean that this background ain't mine? He knows me well and everybody knows him well, we know that he is trustable. What is the problem?

OpenStudy (kc_kennylau):

The rules, in the ultimate purpose, is just to help people :)

OpenStudy (kc_kennylau):

are*

OpenStudy (tkhunny):

@kc_kennylau Helping is, indeed, the goal. Knowing just how to do that is a nontrivial exercise. Generally, this is best accomplished by simply following the rules. It may very well be your background. Thus the "or". A little better background in logic will expose this additional option. I have expressed my position sufficiently.